OpenStudy (anonymous):

please help me understand this problem Quotient rule exercise: x^(30 )/x^10 x^(30 )/x^(–10)

zepdrix (zepdrix):

These are two separate problems, yes?

zepdrix (zepdrix):

For the first one, here is how our Quotient Rule setup would look: \[\Large \left(\frac{x^{30}}{x^{10}}\right)' \quad=\quad \frac{\color{royalblue}{(x^{30})'}(x^{10})-(x^{30})\color{royalblue}{(x^{10})'}}{(x^{10})^2}\] Confused by any of that? We need to take the derivative of the blue parts.

OpenStudy (anonymous):

yes thank you